# On the spectrum of Sturmian Hamiltonians of bounded type in a small coupling regime

### v1

@@ 2.3.2 Dynamical Rectangles
We say that a $C^{1}$ curve contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ that is tangent to $\mathcal{K}^{s}$ connects opposite boundaries of $\tilde{K}^{u}(Q)$ if its endpoints belong to $\partial\tilde{K}_{+}^{u}(Q)$. For any $Q\in\mathbb{T}^{2}$ and points $P,P^{\prime}\in\mathbb{B}_{\frac{1}{10}}(Q)$, we say that a curve $\gamma$ connects $P$ and $P^{\prime}$ if these are the endpoints of $\gamma$, and $\gamma$ is strictly contained in $\overline{B_{\frac{1}{10}}(Q)}$. If $\gamma_{1}$ and $\gamma_{2}$ are disjoint curves contained in $B_{\frac{1}{10}}(Q)$, then $\gamma$ connects these curves if it connects points from these curves.
For each $Q\in\mathbb{T}^{2}$ and cone $K\supset K^{u}$ transversal to $K^{s}$, we say that two $C^{1}$ curves $\gamma_{1}$ and $\gamma_{2}$ contained in $\mathbb{B}_{\frac{1}{10}}(Q)$, tangent to $\mathcal{K}^{s}$, are centered at $Q$ with respect to $\tilde{K}(Q)$, if there is a sub-curve of $\gamma_{1}$ that connects opposite boundaries of $\tilde{K}_{+}(Q)$ and does not intersect $\tilde{K}_{-}(Q)$, and if there is a sub-curve of $\gamma_{2}$ that connects opposite boundaries of $\tilde{K}_{-}(Q)$ and does not intersect $\tilde{K}_{+}(Q)$, or vice versa (see Figure 4).
@@
A rectangle $R\subset B_{\frac{1}{10}}(Q)$ that contains $Q$ in its interior is centered at $Q$ with respect to $\tilde{K}(Q)$ if its stable boundary curves are centered at $Q$ with respect to $\tilde{K}(Q)$ (see Figure 4). We denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$. It should also be noted that ±$\partial^{s}_{\pm}R\cap\tilde{K}^{u}_{\mp}(Q_{i})=\emptyset$.
@@
Figure 4: The right figure depicts two curves $\gamma_{1}$ and $\gamma_{2}$ that are centered at $Q$ with respect to a cone $\tilde{K}(Q)$, and the left figure depicts a dynamical rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$.
Remark 4.
Notice there is a $C^{*}_{0}>0$ such that for each $i=1,\dots,4$, if $\gamma$ is a $C^{1}$ curve tangent to $\mathcal{K}^{s}$ that is contained in $\overline{\mathbb{B}}_{i}$ satisfying $\text{dist}(\gamma,Q_{i})<C^{*}_{0}$, then it is disjoint from $\partial^{s}\mathbb{B}_{i}$ (see Figure 5). Moreover, if $\kappa$ is a curve tangent to $\mathcal{K}^{u}$ with endpoints $P_{1}$ and $P_{2}$ satisfying $P_{1}\in\tilde{K}_{+}(Q_{i})$, $P_{2}\in\tilde{K}_{-}(Q_{i})$, and $\max\left\{\left|P_{1}-Q_{i}\right|,\left|P_{2}-Q_{i}\right|\right\}<C^{*}_{0}$, we must have $\kappa\subset\mathbb{B}_{i}$. That is $\kappa$ is disjoint from $\partial^{u}\mathbb{B}_{i}$.
$Q$$\tilde{K}^{s}(Q)$$\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q)$$\gamma$
$\tilde{K}(Q)$$\tilde{K}^{u}(Q)$$Q$$\kappa$
Figure 5: Illustration of Remark 4. On the top figure, any curve $\gamma$ tangent to $\mathcal{K}^{s}$ that intersects the center “diamond” must stay between the dotted lines. On the bottom figure, any curve $\kappa$ tangent to $\mathcal{K}^{u}$ with endpoints in the center region contained in $\tilde{K}(Q)$, must be contained within the dotted lines. So, one can choose a neighborhood $O$ of $Q$ accordingly.
Remark 5.
If $K\supsetneq K^{u}$ is a cone disjoint from $K^{s}$, then for any rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$, there is a neighborhood $O$ of $Q$ such that for any curve $\gamma$ tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects $O$ must not intersect $\partial^{ns}R$ (see Figure 6). The size of the neighborhood $O$ depends on the cone $K$ and $\text{dist}(\partial R,Q)$.
In particular, if $\{R_{n}\}$ is a family of rectangles that are centered at $Q$ with respect to $\tilde{K}(Q)$ and there is a $l^{*}>0$ such that $\inf_{n}\text{dist}\left(\partial R_{n},Q\right)\geq l^{*}$, then there is a $r=r(K,l^{*})>0$ such that for each $n$ and any curve $\gamma$ tangent to $\mathcal{K}^{u}$, we must have that any component of $\gamma\cap R_{n}$ that intersects $B_{r}(Q)$, must not intersect $\partial^{ns}R_{n}$. That is, for any curve $\gamma$ tangent to $\mathcal{K}^{u}$ connecting $\partial_{+}^{s}R_{n}$ and $\partial_{-}^{s}R_{n}$ with $\gamma\cap B_{r}(Q)\neq\emptyset$, we must have $\gamma\subset R_{n}$ and $\gamma\cap\partial^{ns}R=\emptyset$.
@@
Figure 6: Illustration of Remark 5. The rectangle $R$ is centered at $Q$ with respect to $\tilde{K}(Q)$ and has the property that for any curve $\gamma$ that is tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects the center quadrilateral must stay within the dotted lines and hence not intersect $\partial^{ns}R$.
@@ 3 Main Estimates and Non-Stationary Trace Map Dynamics
For the rest of the paper, let us consider an arbitrary finite set $\mathcal{A}\subset\mathbb{N}$ and an arbitrary sequence $\overline{a}:=(a_{n})$ such that $a_{n}\in\mathcal{A}$ for all $n\in\mathbb{N}$.
Let $M>0$. We will show that there is a $\lambda_{0}>0$ such that for all $\lambda\in[0,\lambda_{0}]$, we have
$\text{dim}_{H}\left(L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}.$ (3.1)
Our approach is to invoke Lemma 5 and choose a neighborhood $O$, that is composed of small neighborhoods containing the points $Q_{i},$ so that the set
$C(\lambda,O):=\{P\in\hat{L}_{\lambda}\setminus O:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\}$
is diffeomorphic to a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}$. We will choose the neighborhood $O$ so that the dimension of this set is bigger than the quantity $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ for small $\lambda$. More specifically, the neighborhood $O$ is chosen in a way so that for a sufficiently small range of $\lambda$, the set $C(\lambda,O)$ contains a Cantor set of thickness larger than $M$.
In the next section, we introduce a series of constants that will be used throughout the proof, and these constants will determine the size of the neighborhoods that make up $O$.
@@ 3.1 Development of Constants
Recalling (2.2), we fix a cone $K:=K_{\beta_{1}}^{u}$, for some $\beta_{1}\in(\beta_{0},0.1)$, so that it strictly contains $K^{u}$ and is transversal to $K^{s}$. We also choose $\beta_{1}$ close enough to $\beta_{0}$ so that the property described in Remark 2 holds for the cone $K$. If $K(x,y)$ denotes the cone $K$ in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, then we denote $\mathcal{K}:=\left\{K(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}$. We set
@@
where the constants $C_{1}$ and $C^{*}_{0}$ are taken from Remarks 3 and 4, respectively. Referring to Lemma 2, we set
@@
$C_{3}:=\sum_{n=-1}^{\infty}\frac{1}{(\mu_{1}-\delta)^{-(n+1)}},$ (3.3)
Remark 6.
There is a $\Theta>0$ such that if $(T_{n})$ is a sequence of maps satisfying Property (C) and $\sup_{n}\|T_{n}\|_{C^{2}}<C_{2}+1$, and if $\gamma$ is a smooth curve tangent to $\mathcal{K}^{u}$ so that $\sum_{i=0}^{n}\text{length}\left(T_{n}\circ\cdots\circ T_{1}(\gamma)\right)<C_{1}C_{3}\frac{\sqrt{2}}{10}$ for each $n$, then $\sum_{i=0}^{n}\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)<\Theta$ for each $n$, where $\text{Ang}(\kappa)$ denotes the supremum of the angles between tangent vectors of $\kappa$.
@@
By Lemma 4 we have
@@
There is a $l^{*}>0$ such that if $L$ is a line through $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then for each $i=1,\dots,4$, there are segments $L^{+}_{i},\ L^{-}_{i}\subset L$ contained in $\mathbb{B}_{i}\setminus\partial^{s}\mathbb{B}_{i}$ such that
@@
where $\theta$ is the angle between boundary lines of $\tilde{K}(Q_{i})$ (or $K$).
Let us choose a $\tilde{n}$ so that $\frac{\left(\mu_{1}-\delta\right)^{\tilde{n}}}{10}>l^{*}$ and set $l^{**}:=\left(\mu_{2}+\delta\right)^{\tilde{n}+1}$. Then, again by Lemma 4, we have
@@
There is a $C^{**}\in\left(0,\frac{1}{4}-\frac{1}{10}\right)$ such that if $L$ is a line through $Q_{0}$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\leq l^{**}$, then
@@
Recalling our constant $C_{1}$ from Remark 3, we will choose $N^{*}\in\mathbb{N}$ such that

### v2

@@ 3.2 Dynamical Rectangles
Definition.
Let $\gamma$ be a $C^{1}$ curve contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ that is tangent to $\mathcal{K}^{s}$. We say that $\gamma$
  • connects opposite boundaries of $\tilde{K}^{u}(Q)$ if its endpoints belong to $\partial\tilde{K}_{+}^{u}(Q)$
  • connects points $P,P^{\prime}\in\mathbb{B}_{\frac{1}{10}}(Q)$, if these are the endpoints of $\gamma$, and $\gamma$ is strictly contained in $\overline{B_{\frac{1}{10}}(Q)}$
  • connects disjoint curves $\gamma_{1},\gamma_{2}\subset B_{\frac{1}{10}}(Q)$ if it connects points of these curves
Definition.
Let $Q\in\mathbb{T}^{2}$ and $K\supset K^{u}$ be a cone transversal to $K^{s}$. We say
  • Two $C^{1}$ curves $\gamma_{1},\gamma_{2}\subset\mathbb{B}_{\frac{1}{10}}(Q)$ tangent to $\mathcal{K}^{s}$, are centered at $Q$ with respect to $\tilde{K}(Q)$, if
    • There is a sub-curve of $\gamma_{1}$ that connects opposite boundaries of $\tilde{K}_{+}(Q)$ and does not intersect $\tilde{K}_{-}(Q)$
    • There is a sub-curve of $\gamma_{2}$ that connects opposite boundaries of $\tilde{K}_{-}(Q)$ and does not intersect $\tilde{K}_{+}(Q)$
    or vice versa (see Figure 6).
  • A rectangle $R\subset B_{\frac{1}{10}}(Q)$ is centered at $Q$ with respect to $\tilde{K}(Q)$ and its stable boundary curves are centered at $Q$ with respect to $\tilde{K}(Q)$. In this case we denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$. (see Figure 6)
Remark 3.
@@
Also, we denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$. It should also be noted that ±$\partial^{s}_{\pm}R\cap\tilde{K}^{u}_{\mp}(Q_{i})=\emptyset$.
3.3 Geometric Properties of Dynamical Rectangles
We now derive two propositions that describe useful geometric properties concerning these local cones and dynamical rectangles. The next proposition is illustrated by Figure 7.
Proposition 1.
There is a $C^{*}_{0}>0$ such that for any $i=1,\dots,4$, the following hold:
  • (1)
    If $\gamma$ is a $C^{1}$ curve tangent to $\mathcal{K}^{s}$ that is contained in $\overline{\mathbb{B}}_{i}$ satisfying
    $\text{dist}(\gamma,Q_{i})<C^{*}_{0},$
    then $\gamma$ is disjoint from $\partial^{s}\mathbb{B}_{i}$.
  • (2)
    If $\kappa$ is a curve tangent to $\mathcal{K}^{u}$ with endpoints $P_{1}$ and $P_{2}$ satisfying $P_{1}\in\tilde{K}_{+}(Q_{i})$, $P_{2}\in\tilde{K}_{-}(Q_{i})$, and
    $\max\left\{\left|P_{1}-Q_{i}\right|,\left|P_{2}-Q_{i}\right|\right\}<C^{*}_{0},$
    we must have $\kappa\subset\mathbb{B}_{i}$. That is $\kappa$ is disjoint from $\partial^{u}\mathbb{B}_{i}$.
Proof.
We prove part (1). Since $\text{dist}\left(\partial^{s}\mathbb{B}_{i},\tilde{K}^{s}(Q_{i})\right)>0$, by construction of $K^{s}$, if $L\subset\overline{\mathbb{B}}_{i}$ is a line segment parallel to a segment of $\partial\tilde{K}^{s}(Q_{i})$, and sufficiently close to $Q_{i}$, we must also have $\text{dist}\left(\partial^{s}\mathbb{B}_{i},L\right)>0.$ It follows that the same is true for any curve $\gamma\subset\overline{\mathbb{B}}_{i}$ that is tangent to $\mathcal{K}^{s}$. The proof of (2) is similar.
@@ 3.3 Geometric Properties of Dynamical Rectangles
Figure 6: The left figure depicts two curves $\gamma_{1}$ and $\gamma_{2}$ that are centered at $Q$ with respect to a cone $\tilde{K}(Q)$, and the right figure depicts a dynamical rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$.
[…]
Figure 7: Illustration of Proposition 1. On the top figure, any curve $\gamma$ tangent to $\mathcal{K}^{s}$ that intersects the center “diamond” must stay between the dotted lines. On the bottom figure, any curve $\kappa$ tangent to $\mathcal{K}^{u}$ with endpoints in the center region contained in $\tilde{K}(Q)$, must be contained within the dotted lines. So, one can choose a neighborhood $O$ of $Q$ accordingly.
The next proposition is illustrated by Figure 8.
Proposition 2.
Suppose $K\supsetneq K^{u}$ is a cone disjoint from $K^{s}$, and $R$ is a rectangle centered at $Q$ with respect to $\tilde{K}(Q)$. Then, there is a neighborhood $O$ of $Q$ such that for any curve $\gamma$ tangent to $\mathcal{K}^{u}$, we must have that any component of $\gamma\cap R$ that intersects $O$ must not intersect $\partial^{ns}R$.
Proof.
The proof is analogous to that of Proposition 1. ∎
It should be noted that the size of the neighborhood $O$ in this proposition depends only on the cone $K$ and $\text{dist}(\partial R,Q)$.
@@
Figure 8: Illustration of Proposition 2. The rectangle $R$ is centered at $Q$ with respect to $\tilde{K}(Q)$ and has the property that for any curve $\gamma$ that is tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects the center quadrilateral must stay within the dotted lines and hence not intersect $\partial^{ns}R$.
@@ 4 Main Estimates and Non-Stationary Trace Map Dynamics
We will now translate Theorem 3 into a problem concerning a sequence of Anosov maps of $\mathbb{T}^{2}$. For the rest of this section, let us consider an arbitrary finite set $\mathcal{A}\subset\mathbb{N}$ and an arbitrary sequence $\overline{a}:=(a_{n})$ such that $a_{n}\in\mathcal{A}$ for all $n\in\mathbb{N}$.
Recalling Lemma 5, given $\rho>0$, we set $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$ to be the curve given by $F_{\lambda}^{-1}\left(L_{\lambda}\setminus O_{\rho}\right)$. Now, for appropriate range of $\lambda$, we define
$C(\lambda,O):=\{P\in\hat{L}_{\lambda}\setminus O:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\}.$
Theorem 3 will follow from
Theorem 4.
For each $M>0$, there is a $\rho>0$ and $\lambda_{0}\in\left(0,\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}\right)$, such that
$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (4.1)
for all $\lambda\in[0,\lambda_{0}]$.
The proof of this theorem will be developed throughout Section 4 and will be concluded in Subsection 4.8. Let us fix $M>0$.
4.1 Outline of Proof of Theorem 4
We outline the proof of the argument to come
  • In Subsection 4.2, we introduce a series of constants that will be used throughout the proof of this theorem, and these constants will determine $\rho$, and hence the size of the neighborhoods that make up $\tilde{O}_{r(\rho)}$.
  • Subsection 4.3 is dedicated to choosing $\rho$ explicitly, and once chosen, in Subsection 4.4, we extend each map $\tilde{T}_{\lambda,a}$, for a small range of $\lambda$ and $a\in\mathcal{A}$, to a toral Anosov map, and these maps will satisfy a common cone condition.
  • In Subsection 4.5, we then derive useful dynamical properties of these maps such as a bounded distortion estimate, and a non-stationary version of the Stable Manifold Theorem for these maps.
  • In Subsection 4.6, we construct families of sequences of dynamical rectangles containing each component of $\tilde{O}_{r(\rho)}$, that will be useful in controlling the dynamics of our maps near this open set. The stable sides of these rectangles will be composed of the stable manifolds constructed in Subsection 4.5.
  • Lastly, in Subsection 4.7, for each $\lambda$ sufficiently near 00, we construct a subset of $\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)$, that is roughly composed of the points whose orbits do not enter the interior of the constructed rectangles from Subsection 4.6. The proof is concluded in Subsection 4.8 by showing that this subset has large thickness (see Subsection 4.7.1 for a formal definition), and in particular, Hausdorff dimension greater than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
@@ 4.2 Development of Constants
Recalling (2.2), we fix a cone $K:=K_{\beta_{1}}^{u}$, for some $\beta_{1}\in(\beta_{0},0.1)$, so that it strictly contains $K^{u}$ and is transversal to $K^{s}$. We also choose $\beta_{1}$ close enough to $\beta_{0}$ so that the property described in Lemma 6 holds for the cone $K$. If $K(x,y)$ denotes the cone $K$ in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, then we denote $\mathcal{K}:=\left\{K(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}$.
We set
@@
where the constants $C_{1}$ and $C^{*}_{0}$ are taken from Lemma 7 and Proposition 1, respectively. Referring to Lemma 2, we set
@@
$C_{3}:=\sum_{n=-1}^{\infty}\frac{1}{(\mu_{1}-\delta)^{-(n+1)}}.$ (4.3)
Lemma 8.
There is a $\Theta>0$ such that the following holds: If $(T_{n})$ is a sequence of maps satisfying Property (C) and $\sup_{n}\|T_{n}\|_{C^{2}}<C_{2}+1$, and if $\gamma$ is a smooth curve tangent to $\mathcal{K}^{u}$ so that
$\sum_{i=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}\text{length}\left(T_{i}\circ\cdots\circ T_{1}(\gamma)\right)<C_{1}C_{3}\frac{\sqrt{2}}{10}$
for some $m\geq 0$, then
$\sum_{i=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)<\Theta,$
where $\text{Ang}(\kappa)$ denotes the supremum of the angles between tangent vectors of $\kappa$.
Proof.
Let $\gamma$ be a curve tangent to $\mathcal{K}^{u}$. From Property (C) and since the $C^{2}$ norms of each $T_{n}$ are uniformly bounded, the curvature of each curve $T_{i}\circ\dots\circ T_{1}(\gamma)$, $1\leq i\leq m$, is uniformly bounded. Hence, there is a constant $\Theta^{\prime}>0$, that only depends on $C_{2}$ and the cones and constants in Property (C), such that
$\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)\leq\Theta^{\prime}\text{length}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)$
for each $1\leq i\leq m$. The lemma follows.
@@
Let us set $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$. By Lemma 4 and Proposition 1 we have
@@
There is a $l^{*}>0$ such that if $L$ is a line through $Q_{0}$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then for each $i=1,\dots,4$, there are segments $L^{+}_{i},\ L^{-}_{i}\subset L$ contained in $\mathbb{B}_{i}\setminus\partial^{s}\mathbb{B}_{i}$ such that
@@
where $\theta$ is the angle between boundary lines of $\tilde{K}(Q_{i})$ (explicitly, $\theta=\theta_{0}(\beta_{1})$, see discussion before (2.2)).
Proof.
Let $i\in\{1,2,3,4\}$. By Lemma 4, we can choose a $l^{*}>0$ such that if $L$ is a line through $Q_{0}$ of slope $\frac{-1}{[b_{1},b_{2},\dots]}$, $b_{i}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then there is a segment $L^{\prime}_{i}\subset L$ such that
  • $L^{\prime}_{i}\subset\tilde{K}_{+}(Q_{i})$ and connects the boundaries of $\tilde{K}_{+}(Q_{i})$
  • For each $P\in L_{i}^{\prime}$, we have $|P-Q_{i}|\leq\min\left\{C_{0}^{*},\ \cos(\theta)\frac{C^{*}}{4}\right\}.$
By increasing $l^{*}$ if needed, we can even assume that $L^{\prime}_{i}$ connects the boundary lines of $\tilde{K}_{+}(Q_{i})$.
Let $L$ be such a line and let $L^{+}_{i}$ be the longest segment of $L\cap\overline{\mathbb{B}}_{i}$ that contains $L^{\prime}_{i}$. Then, by Proposition 1, since
$\text{dist}\left(L_{i}^{+},Q_{i}\right)<C_{0}^{*},$
we have that $L_{i}^{+}\cap\partial^{s}\mathbb{B}_{i}=\emptyset.$ By increasing $l^{*}$ again if needed, we can also ensure that $L_{i}^{+}$ connects the components of $\partial^{u}\mathbb{B}_{i}$ and Proposition 1 will still guarantee that this extended segment is disjoint from $\partial^{s}\mathbb{B}_{i}$. By repeating this argument for each $i$ and with $+$ replaced with --, the lemma is proved.
Let us choose a $\tilde{n}$ so that $\frac{\left(\mu_{1}-\delta\right)^{\tilde{n}}}{10}>l^{*}$ and set $l^{**}:=\left(\mu_{2}+\delta\right)^{\tilde{n}+1}$. For $m\geq 1$, denote
$\text{Orb}_{0}\left(Q_{0}\right):=\left\{\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(Q_{0}\right)\right\}_{n\geq 0}\cup\left\{Q_{0}\right\}.$
It is a calculation to check that $\text{Orb}_{0}\left(Q_{0}\right):=\bigcup_{m\geq 0}\text{Orb}_{0}\left(Q_{0}\right)$ is a finite set and has distance from each $Q_{i}$ that is at least $\frac{1}{4}$. Again by Lemma 4, we have
@@
There is a $C^{**}\in\left(0,\frac{1}{4}-\frac{1}{10}\right)$ such that if $L$ is a line through $Q$, for some $Q\in\text{Orb}_{0}\left(Q_{0}\right)$, of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\leq l^{**}$, then
@@
Proof.
The proof is analogous to that of Lemma 9 but through an application Lemma 4 part (2). The quantity $\frac{1}{4}-\frac{1}{10}$ comes from the distance between $\text{Orb}_{0}\left(Q_{0}\right)$ and each $Q_{i}$, and the size of each $\mathbb{B}_{i}$.
Recalling our constant $C_{1}$ from Lemma 7, we will choose $N^{*}\in\mathbb{N}$ such that